rods»view/108

Status Summary for ROD ID 108

tickets

<<Prev | Next>>

Test NameStatusNotesModified
Board Status 0 Replaced QPLL Crystal X3, now locks to TTC and internal. -AC 5/2/08 Still high, but works fine. -AC 4/30 5V current is too high (0.68 amps). During initial testing loop data test failed. With CGs new mig and ise 10.1 both top and bottom pass loop data (27/05/08 alex). After a reset of the TTC this ROD takes ~5 - 10 minutes to regain QPLL. Adam has possibly traced this to a fault on th 2.5 VCC line. April,2015: R11 (for 2.5V power line at the bottom) was a white color which looked different from the R11 of any other board, so the component was replaced. ZG2015April29: No VME connection to the board. J5 chain does not work. Loading VME via J4 does not help to bring life to VME. 29-04-2015
fuses 1
17-06-2008
U31 1
17-06-2008
U32 1
17-06-2008
U33 1
17-06-2008
U34 1
17-06-2008
U35 1
17-06-2008
U81 1
17-06-2008
U82 1
17-06-2008
I5V_initial 1
17-06-2008
I3V initial 1
17-06-2008
proms load 1
17-06-2008
I5V proms 1
17-06-2008
I3V proms 1
17-06-2008
V4T temp 1
17-06-2008
V4B temp 1
17-06-2008
connectivity 1
17-06-2008
loop data 1
All bits shifted by 1 and top byte is 00. -AC 5/2/08. see sttlog108 for failed fifo test 5/14 BM. With CGs new mig and ise 10.1 top and bottom pass, although the top has only been tested with 30k evts (27/05/08 alex).
17-06-2008
slink 1
17-06-2008
dram 1
17-06-2008
nse communication 1
17-06-2008
sram
17-06-2008
optical loop 1
17-06-2008
patch panel 1
17-06-2008
qpll lock 1
17-06-2008
ttc connectivity 1
17-06-2008
test fifo 1
17-06-2008
ddr reset
17-06-2008
assembly 1
17-06-2008